What mass of aluminum hydroxide is formed from 12gm of aluminum carbide which 144 gm

Grade:10

1 Answers

Arun
5 years ago
According to balance equation 
Al4C3 + 12 H2O → 3CH4 + 4Al(OH)3
 
144 gm.                                           312 gm

144 gm gives 312 gm aluminium hydroxide
 
hence 1 gm will give =312/144
 
hence 12 gm will give = 312 * 12 / 144 = 26 gm